2.1 Suppose P(A) = 0.75 and P(B) = 0.82. Find the following:


2.1.1 P (A and B) if A and B are independent. (2)


2.1.2 P (not A and not B), if A and B are independent. (2)


2.1.3 P (A and B), if A and B are mutually exclusive. (1)


2.2 Solve the following inequality. (3)


x2 - 5x + 6 ≥ 0


1
Expert's answer
2022-05-31T12:15:45-0400

2.1.1


"P(A\cap B)=P(A)P(B)"

"=0.75(0.82)=0.615"

2.1.2


"P(A^C\cap B^C)=P(A^C)P(B^C)"

"=(1-0.75)(1-0.82)=0.045"

2.1.3


"P(A\cap B)=0"

2.2


"x^2 - 5x + 6 \u2265 0"

"(x-2)(x-3)\ge0"

"x\in(-\infin, 2]\cup [3, \infin)"